Technician team had identified that a Power injector was dead and had replaced it. services are restored now.
Posted Jul 29, 2025 - 12:16 AEST
Update
The fiber team has now been dispatched to the site for further troubleshooting
Posted Jul 29, 2025 - 09:34 AEST
Update
Suspect that a third-party contractor who was working on-site may have damaged the fiber, resulting in the outage. The fiber team will be dispatched tomorrow morning (29th) to conduct further investigation.
Posted Jul 28, 2025 - 20:18 AEST
Update
DGTEK Technician is at site and further investigating the issue.
Posted Jul 28, 2025 - 17:47 AEST
Update
DGTEK Field Technician is dispatched to the site to investigate further.
Further updates will be given as they become available.
Posted Jul 28, 2025 - 16:19 AEST
Investigating
Engineers have observed network alarms at the listed site which are causing disruption to some customer services. Engineers are now investigating the issue. Further updates will be given as soon as they are available.
We apologise for any inconvenience caused and are working to resolve issue as soon as possible.